Rahul has a wonderful machine. If you put any number in that machine, it first multiply the number by 4 then subtract 8 from the result and then add the successor of the number inserted to the result. If Martin puts the greatest three digit number into the machine, what will be the result?  

Grade:10

3 Answers

Swoyana Das
3 years ago
so the greatest three digit no is 999
then when we insert it the machine multiplies the number with 4
so 999*4=3996
then it subtracts 8 from the result
so 3996-8=3988
atq it adds the successor of the number inserted to the result
so the successor of 999 id 1000
so 3988+1000=4988
here you go with the answer=4988
Sarthak Shahade
one year ago
Rahul has a wonderful machine. If you put any number in that machine, it first multiply the number by 4 then subtract 8 from the result and then add the successor of the number inserted to the result. If Martin puts the greatest three digit number into the machine, what will be the result? 
Sarthak Shahade
one year ago
From the numbers given in the following figure, Anika multiplies the smallest prime number and the greatest number and then add the remaining numbers to the product. The resulting number is :